OpenStudy (anonymous):

The point 5/2,y is a solution to the equation 2x+5y+30=0. What is the value of y

OpenStudy (anonymous):

welcome to open study please give me a second

OpenStudy (anonymous):

what do you think it is?

OpenStudy (michele_laino):

hint: please you have to replace \(x=5/2\) into the eqaution, so we can write: \[\Large 2 \cdot \frac{5}{2} + 5y + 30 = 0\] please solve for y